OpenStudy (across):

Suppose that \(f\) is some polynomial of prime order \(p\) over the rationals and that it has exactly two nonreal roots. I just need one idea to show that the Galois group of its splitting field is isomorphic to the symmetric group of \(p\) letters.

OpenStudy (across):

I've so far been able to show that \(x\mapsto\overline x\) is in \(\text{Gal}\left(K/\mathbb Q\right)\), where \(K\) is the splitting field of \(f\), and has order \(2\). Also, if \(\alpha\) is one of those nonreal roots, then \(\left[\mathbb Q\left(\alpha\right):\mathbb Q\right]=p\), which implies that \(\text{Gal}\left(K/\mathbb Q\right)\) has an element of order \(p\) by Cauchy's theorem.

OpenStudy (across):

I may have read that if \(\alpha\) is a \(2\)-cycle and \(\beta\) is a \(p\)-cycle (prime), then \(\langle\alpha,\beta\rangle\) generate the symmetric group of \(p\) letters. Is this even true?

OpenStudy (across):

I think I just need to show that a \(2\)-cycle and a \(p\)-cycle, where \(p\) is prime, generate \(S_p\), if anyone might want to help me with that.

OpenStudy (across):

@LetsLearn2000, my name is Susanne. Pleasure to meet you.

OpenStudy (anonymous):

Im Vanessa. Nice to meet you as well Susanne.

OpenStudy (across):

I've been a member of this site for five years. It's all new faces now. T_T All of my friends are gone!

OpenStudy (anonymous):

yeah, its a lot of people just like you who have been on here for years, and some people are new.

OpenStudy (across):

I don't recognize half of the guys here.

OpenStudy (anonymous):

yeah, its surely some new faces lol. Could you continue helping out with those questions?

OpenStudy (thomas5267):

@ganeshie8 @Kainui Hardcore stuff that is usually not seen on here.

OpenStudy (kainui):

Nope, haven't taught myself finite fields yet, it's on my todo list though.

OpenStudy (bobo-i-bo):

It's out of my knowledge as well... but I do know that every group of prime order is cyclic... if that helps at all :P

OpenStudy (thomas5267):

What I can gather from the internet is that a 2-cycle that swaps adjacent elements and a n-cycle together generates the symmetric group of order n.

OpenStudy (thomas5267):

*symmetric group of n letters, not order n. I am not familiar with group theory at all.

OpenStudy (across):

@thomas5267, so the \(n\)-cycle need not be prime. Hmm. I'm still struggling to prove this.

OpenStudy (across):

If I prove this, I will be able to say that \(S_p\cong\langle\tau,\sigma\rangle=\text{Gal}\left(K/\mathbb Q\right)\) since \(\left[K:\mathbb Q\right]\) can be at most \(p!\), the order of \(S_p\).

OpenStudy (thomas5267):

What is that 2-cycle in question? Did you prove a 2-cycle exists or were you able to explicitly construct such 2-cycle?

OpenStudy (thomas5267):

I suspect since the symmetric group has p letters, it might be the case that any 2-cycle is good enough?

OpenStudy (across):

It's the first sentence of my first post.

OpenStudy (across):

Never mind. I got it. I reduced the problem to showing that the set of transpositions of \(S_n\) of the form \(\left(i,i+1\right)\), where \(0\leqslant i< n-1\), generates \(S_n\). Then I will use the fact that \[\left(1,2,\dots,n\right)^i\left(1,2\right)\left(1,2,\dots,n\right)^{-i}=\left(i+1,i+2\right).\]